OpenStudy (anonymous):

How to solve sin2x-sinx=0? Thanks

OpenStudy (anonymous):

is it \[\sin^2(x)-\sin(x)=0\] or \[\sin(2x)-\sin(x)=0\]?

OpenStudy (anonymous):